# GazeNoter: Co-Piloted AR Note-Taking via Gaze Selection of LLM Suggestions to Match Users’ Intentions

### v1

@@ 4.2. Task and Procedure
The order of the methods was counterbalanced. Besides one video for the training, one of the other five videos was randomly played in a method. The time and steps of each note and each feature of GazeNoter, and the number of notes, keywords and sentences users selected were recorded. Furthermore, participants needed to fill out a questionnaire with a 7-point Likert scale. 11 questions were for each method. 4 of them, including intention, reminding, quality and inspiration, were for notes from each method, and 4 were based on NASA-TLX, as shown in Figure 7. In quality, it only represents the quality of the notes but was irrelevant to whether the notes matched users’ intentions, which was rated in intention. In reminding, it means whether the notes from a method could remind users of their thoughts or opinions. The 4 questions for notes were used to evaluate not only the overall ratings but also each note for deeper investigation. Notably, since no user input was in (A), and the additional comparison was only between (A) and (G), participants only scored the 4 questions for the notes in (A). They were interviewed for some feedback. The study lasted approximately two hours, including the introduction, training sessions, breaks between methods and interview.
@@ 4.3. Result and Discussion
The results are shown in Figure 6 to 7. Repeated measures ANOVA and Bonferroni correction for post-hoc pairwise tests were used to analyze the objective, quantitative data, including time and numbers of notes and keywords. A Friedman test and Wilcoxon signed-rank tests with Bonferroni correction for post-hoc pairwise tests were utilized to analyze the results of the subjective questionnaire.
@@ 4.3.1. Quantitative Results
The time of a note was measured from the first selection of a keyword to the last selection of recording the note in (S) and (G), and from typing the first letter to pressing the last “return” button to complete the note in (T). The number of keywords per note was calculated by including notes with only keywords and excluding notes with complete sentences. Furthermore, the percentage of notes with complete sentences is 28% in (T), 73% in (S), and 78% in (G), and the others are notes with only keywords. The percentage of quick notes is 21% in (S) and 18% in (G). Quick notes were the notes recorded with only context (or also customized) keywords and completed with less than 2s in the final recording step via double-click, which was less than the average candidate sentence generation time of 2.89s. This condition indicates that participants intended to take notes quickly by selecting keywords without considering reading candidate sentences, which means only (a) (b) or (a) (b) (c) steps in Figure 2. The average time of a quick note in (S) is 5.22s for context keywords only and 6.59s for customized keywords included. The average time of a quick note in (G) is 5.05s for context keywords only and 5.85s for customized keywords included. These demonstrate how fast the proposed system could be if quick notes are required. The percentage of beyond-context notes, meaning notes with derivative keywords selected, is 22% in (S) and 28% in (G). This shows the importance of beyond-context notes in note-taking.
@@
4.3.2. Qualitative Results
@@ 4.3.3. Additional Discussions
In the objective and quantitative results, P1, P4, P7, P8, and P10 commented that context keywords were easy to access and organized, which encouraged them to take more notes. P1, P7 and P9 appreciated the LLM-generated sentence feature and were more willing to take notes due to no need to manually type. For the time of a note and number of keywords per note, all participants reported that the AR content did not disrupt their viewing of the speech, except when momentarily using the derivative keyword feature, which required minimal time. 7 participants (P3, P5, P7, P8, P9, P10, P11) mentioned that they were accustomed to keeping the AR display on throughout the entire speech to show context keywords, which facilitated quicker and easier keyword selection. P7 elaborated that “Always displaying the AR content allowed me to select the context keywords and sentences as soon as they were extracted and generated, respectively. However, achieving this on the smartphone (S) required constantly checking the screen, which was very distracting and hindered me from performing more keyword and sentence selection.” Furthermore, P3, P7, P10 and P11 highlighted the seamless transition between engaging with the speech and note-taking on the AR headset in (G). P10 mentioned that “I could take my time to take notes, switch my focus from the AR content to the speech for a while to stay engaged, and then switch back for note-taking.” The seamless transition in (G) enhanced the flexibility of note-taking and seemed to increase the time of a note due to the inclusion of time spent on engaging with the speech. For (T), P3, P11 and P12 mentioned recording minimal keywords due to the slow typing speed, resulting in a shorter note-taking time. These could contribute that the time in (S) and (G) are similar but the number of keywords in (G) is significantly higher than the others.
@@
For the 4 questions for notes, all participants commented that upon selecting desired keywords, at least one of the generated candidate sentences would match their intentions. This observation could be supported by the results that the significantly higher number of keywords per note in (G) leads to its intention score also being significantly higher than the others. For within-context notes, most participants except for P6 and P8 could select desired context and customized keywords. The average intention score from each within-context note, as rated by participants, is 6.5. However, P6 and P8 encountered a situation where the desired keywords were heard but not extracted as context keywords, diminishing the relevance between the notes and their intentions. For beyond-context notes, half participants (P1, P3, P4, P5, P7, P9) successfully selected desired keywords, subsequently recording an adequate candidate sentence as a note. Nevertheless, the others struggled to always select desired derivative keywords, resulting in less relevant notes. This results in a lower average intention score of 6 from each beyond-context note.
@@
Regarding distraction and cognitive load, all participants agreed that maintaining visual attention on the ongoing speech while taking notes was a clear advantage of using the AR headset in (G). This reduced distraction and cognitive load caused by missing out on the speech content. However, P5 and P8 mentioned that the learning curve for gaze selection could initially increase their cognitive load, but it would become highly efficient and convenient once they were familiar with it. For (T), all participants reported that the slow typing speed and typos led to high cognitive load and even frustration. P1, P6 and P10 further commented that the need to structure the notes, even with minimal keywords, greatly increased their cognitive load. These contribute to significantly higher distraction, cognitive load, and frustration in (T). In terms of subtlety, 6 participants (P2, P4, P9, P10, P11, P12) agreed that (G) only requiring minimal eye and finger movement made it imperceptible. On the contrary, 7 participants (P1, P2, P4, P6, P10, P11, P12) pointed out that looking down at the smartphone was quite noticeable and tiring in both (T) and (S), especially in (T) due to the longer duration for manual typing. These result in (G) being the most subtle method with the least physical effort. For social acceptance, most participants commented that taking notes during a speech was common and would not be considered impolite, even if constantly looking at the smartphone. This causes no statistical significance in social acceptance. For usability, all participants appreciated the context keyword feature in (G) as automatic highlights of the speeches, which was not achieved in (S) due to the need to look down the smartphone. Furthermore, P4, P8, P9 and P10 found the selection feature in (S) and (G) instead of manual typing to be very useful. P10 elaborated that “This combination of AR and AI overcomes all the shortcomings of real-time note-taking.” These result in significantly better usability in (G). For note refinement, only P2 and P10 selected two notes recorded with only keywords to review the transcripts but did not modify or refine the notes, respectively.
@@ 5. User Study 2: Walking Meeting
In the previous study, the performance of GazeNoter in the speech scenario was evaluated in a static condition with less interference. We further intended to evaluate its performance in a dynamic condition, so attending walking meetings was the scenario examined in this study. Moreover, walking meetings involve interactive communication and occasional face-to-face discussion, resulting in frequent eye contact and less time for note-taking. Hence, the mental pressure in this scenario could be higher compared to simply listening to speeches in the previous study.
@@ 5.1. Setup, Task and Procedure
The setup was similar to the previous study. 12 participants (6 female) aged 22-27 (mean: 25.08), who had not participated in the prior study, were recruited for this study. Instead of sitting on a chair, an introducer and a participant walked around a room measuring 3.5m $\times$ 4m for walking meetings. The introducer initiated a topic and spent 3 to 5 minutes describing it, followed by a discussion with the participant during walking meetings. Each walking meeting was about 10 minutes. The participant could freely take notes throughout the description and discussion. However, the participant might need to review the notes for discussion with the introducer. This increased the mental pressure and required more urgent real-time note-taking. As in the previous study, three methods (T), (S) and (G) were examined in each walking meeting, and an additional comparison between (A) and (G) was performed. Similarly, the order of the methods was counterbalanced, and participants filled out a questionnaire and were interviewed after the experiment. The study took about two hours.
@@ 5.2. Results and Discussion
The results are shown in Figure 9 to 10. Repeated measures ANOVA and Bonferroni correction for post-hoc pairwise tests were used to analyze the objective, quantitative data. A Friedman test and Wilcoxon signed-rank tests with Bonferroni correction for post-hoc pairwise tests were utilized to analyze the subjective questionnaire.
@@ 5.2.1. Quantitative Results
For the objective and quantitative results, significant main effects are revealed in number of notes ($F_{2,22}=9.81,p<0.01$), time of a note ($F_{2,22}=26,p<0.001$) and number of keywords per note ($F_{2,22}=16.97,p<0.001$). Post-hoc pairwise tests show significant differences between (T, G) and (S, G) in the number of notes and time of a note, and among all pairs in the number of keywords per note. The percentage of notes with complete sentences is 23% in (T), 37% in (S), and 63% in (G). The quick note percentage is 42% in (S) and 26% in (G). The average time of a quick note in (S) is 5.96s for context keywords only and 7.45s for customized keywords included. The average time of a quick note in (G) is 4.76s for context keywords only and 6.26s for customized keywords included. This scenario leads participants to spend less time on note-taking and prefer quick notes compared to the previous study. For the results of the subjective questionnaire, The beyond-context note percentage is 14% in (S) and 18% in (G). Significant main effects are revealed in intention ($\chi^{2}(2)=22.14,p<0.001$), reminding ($\chi^{2}(2)=15.95,p<0.001$), quality ($\chi^{2}(2)=18.57,p<0.001$), inspiration ($\chi^{2}(2)=13.35,p<0.01$), distraction ($\chi^{2}(2)=22.37,p<0.001$), cognitive load ($\chi^{2}(2)=14.63,p<0.01$), frustration ($\chi^{2}(2)=14.00,p<0.01$), physical effort ($\chi^{2}(2)=17.15,p<0.001$), subtlety ($\chi^{2}(2)=22.14,p<0.001$), social acceptance ($\chi^{2}(2)=17.66,p<0.01$), and usability ($\chi^{2}(2)=21.14,p<0.001$). Post-hoc pairwise tests show significant differences among all pairs in all factors, except between (S, T) in social acceptance. The differences among the methods are more apparent in this scenario. For the additional comparison between (A) and (G), significant main effects are revealed in intention ($\chi^{2}(1)=12.00,p<0.01$) and reminding ($\chi^{2}(1)=12.00,p<0.01$), and no significant main effects are found in quality ($\chi^{2}(1)=0.14,p=0.70$) and inspiration ($\chi^{2}(1)=2.67,p=0.10$), which is consistent with the previous study.
@@
5.2.2. Qualitative Results
Similar feedback and comments were mentioned for several factors as in the previous. We focus on discussing the differences caused by the dynamic condition, interactive communication, and occasional face-to-face discussion in the walking meeting scenario. (G) significantly outperforms both (T) and (S) in the number of notes and time of a note. 6 participants (P2, P4, P9, P10, P11, P12) specified that the diverted attention caused by looking down at the phone in (T) and (S) during discussions led them to take fewer notes and spend less time on each note. P10 mentioned the necessity of seeking permission to take notes in (S) and (T) to avoid being perceived as impolite. P7 pointed out that they had to pause the discussion and sometimes even stop walking to take notes in (S) and (T), which hindered the ongoing conversation. As for (G), two users (P2 and P4) perceived the necessity to slightly move their eyes for interaction during face-to-face conversation diminishing social acceptance while others found it natural to glance away occasionally, enabling them to take notes. These also result in significantly higher social acceptance in (G) compared to (T) and (S), and underscore the influence of social settings on note-taking during occasional face-to-face discussions in walking meetings. Regarding subtlety and physical effort, P10, P10, P10, and P10) specified that constantly switching attention among the walking path, the phone, and the introducer with frequent head movement in (T) and (S) was physically demanding and less subtle. Thus, (G) significantly outperforms (T) and (S) in physical effort and subtlety. In addition, the longer duration for manual typing in (T) results in (T) being significantly less subtle than (S) in walking meetings.

### v2

@@ 4.2.3. Measures
The order of the methods was counterbalanced. Besides one video for the training, one of the other five videos was randomly played in a method. The time and steps of each note and each feature of GazeNoter, and the number of notes, keywords and sentences users selected were recorded. Furthermore, participants needed to fill out a questionnaire with a 7-point Likert scale. 11 questions were for each method. 4 of them, including intention, reminding, quality and inspiration, were for notes from each method, and 4 were based on NASA-TLX, including cognitive load, frustration, physical effort and usability, as shown in Figure 8. The other 3, including distraction, subtlety and social acceptance, were specifically added to understand participants’ experiences when using these three methods to perform note-taking. In quality, it only represents the quality of the notes but was irrelevant to whether the notes matched users’ intentions, which was rated in intention. In reminding, it means whether the notes from a method could remind users of their thoughts or opinions. The 4 questions for notes were used to evaluate not only the overall ratings but also each note for deeper investigation. Notably, since no user input was in (A), and the additional comparison was only between (A) and (G), participants only scored the 4 questions for the notes in (A). They were interviewed for some feedback. The study lasted approximately two hours, including the introduction, training sessions, breaks between methods and interview.
@@ 4.3. Result and Discussion
The results are shown in Figure 7 to 8. Repeated measures ANOVA and Bonferroni correction for post-hoc pairwise tests were used to analyze the objective, quantitative data, including time and numbers of notes and keywords. A Friedman test and Wilcoxon signed-rank tests with Bonferroni correction for post-hoc pairwise tests were utilized to analyze the results of the subjective questionnaire.
@@ 4.3.1. Quantitative Results
The time of a note was measured from the first selection of a keyword to the last selection of recording the note in (S) and (G), and from typing the first letter to pressing the last “return” button to complete the note in (T). The number of keywords per note was calculated by including notes with only keywords and excluding notes with complete sentences. Furthermore, the percentage of notes with complete sentences is 28% in (T), 73% in (S), and 78% in (G), and the others are notes with only keywords. The percentage of quick notes is 21% in (S) and 18% in (G). Quick notes were the notes recorded with only context (or also customized) keywords and completed with less than 2s in the final recording step via double-click, which was less than the average candidate sentence generation time of 2.89s. This condition indicates that participants intended to take notes quickly by selecting keywords without considering reading candidate sentences, which means only (a) (b) or (a) (b) (c) steps in Figure 3. The average time of a quick note in (S) is 5.22s for context keywords only and 6.59s for customized keywords included. The average time of a quick note in (G) is 5.05s for context keywords only and 5.85s for customized keywords included. These demonstrate how fast the proposed system could be if quick notes are required. Furthermore, by comparing steps and time of (G) in Figure 13 and 14, we obtain that 4 seconds could allow the selection of about 2 to 4 context keywords in a note. The percentage of beyond-context notes, meaning notes with derivative keywords selected, is 22% in (S) and 28% in (G) (Figure 12). This shows the importance of beyond-context notes in note-taking. Furthermore, since derivative keywords in (G) could cause more occlusion on the AR headset, it is noteworthy that derivative keywords are displayed for only 27% of the total note-taking time. The percentage during the entire speech duration is even lower, which indicates that distraction is not severe.
@@
Study 1 quantitative results of the three methods, (T), (S) and (G). Showing (G) outperforms (S) and (T) in note quantitative and quality. The dashed line represents the qualitative results of the additional comparison between (A) and (G), illustrating that note outcome of (G) excels in terms of matching user intention and serves as a more effective reminder.
4.3.2. Subjective Scale Results
@@
Study 1 qualitative and subjective scale of the user study on a 7-point Likert scale. The dashed line on the left highlights the 4 questions for notes from each method. The dashed line on the right highlights the scales that lower scores indicate better performance. Results show that (G) significantly outperforms (S) and (T) in various aspects, including intention, distraction, cognitive load, and usability.
@@ 4.3.3. Additional Discussions
In the objective and quantitative results, P1, P4, P7, P8, and P10 commented that context keywords were easy to access and organized, which encouraged them to take more notes. P1, P7 and P9 appreciated the LLM-generated sentence feature and were more willing to take notes due to no need to manually type. For the time of a note and number of keywords per note, all participants reported that the AR content did not disrupt their viewing of the speech, except when momentarily using the derivative keyword feature, which required minimal time. 7 participants (P3, P5, P7, P8, P9, P10, P11) mentioned that they were accustomed to keeping the AR display on throughout the entire speech to show context keywords, which facilitated quicker and easier keyword selection. P7 elaborated that “Always displaying the AR content allowed me to select the context keywords and sentences as soon as they were extracted and generated, respectively. However, achieving this on the smartphone (S) required constantly checking the screen, which was very distracting and hindered me from performing more keyword and sentence selection.” This is consistent with the findings in the note composition distribution in Figure 12 of the appendix, where half of the notes in (S) consist solely of context keywords, requiring minimal time and steps.
Furthermore, P3, P7, P10 and P11 highlighted the seamless transition between engaging with the speech and note-taking on the AR headset in (G). P10 mentioned that “I could take my time to take notes, switch my focus from the AR content to the speech for a while to stay engaged, and then switch back for note-taking.” The seamless transition in (G) enhanced the flexibility of note-taking and seemed to increase the time of a note due to the inclusion of time spent engaging with the speech. For (T), P3, P11 and P12 mentioned recording minimal keywords due to the slow typing speed, resulting in a shorter note-taking time. These could contribute that the time in (S) and (G) are similar but the number of keywords in (G) is significantly higher than the others. This outcome is also supported by the data presented in Figure 13 and 14 in the appendix, where the time distribution for (S) and (G) is similar but more steps are in (G), indicating more efficient note-taking.
4 Questions for Notes
@@
All participants commented that upon selecting desired keywords, at least one of the generated candidate sentences would match their intentions. This observation could be supported by the results that the significantly higher number of keywords per note in (G) leads to its intention score also being significantly higher than the others. For within-context notes, most participants except for P6 and P8 could select desired context and customized keywords. The average intention score from each within-context note, as rated by participants, is 6.5. However, P6 and P8 encountered a situation where the desired keywords were heard but not extracted as context keywords, diminishing the relevance between the notes and their intentions. For beyond-context notes, half participants (P1, P3, P4, P5, P7, P9) successfully selected desired keywords, subsequently recording an adequate candidate sentence as a note. Nevertheless, the others struggled to always select desired derivative keywords, resulting in less relevant notes. As a result, although the average score of each note on intention for beyond-context notes remains high at around 6 points, it is still slightly lower than that for within-context notes (Figure 15). For the additional subjective scales, the 4 questions for each note in Figure 15, statistical analysis was not performed since the number of within-context and beyond-context notes are different. The scores of beyond-context notes are similar to or slightly lower than those of within-context notes. This indicates that participants were still generally satisfied with the beyond-context notes.
@@
Subjective Task Loads
Regarding distraction and cognitive load, all participants agreed that maintaining visual attention on the ongoing speech while taking notes was a clear advantage of using the AR headset in (G). This reduced distraction and cognitive load caused by missing out on the speech content. P7 and P10 stated that they usually read all candidate sentences in (G) to create more complete and longer notes, which could even inspire new ideas. However, P5 and P8 mentioned that the learning curve for gaze selection could initially increase their cognitive load, but it would become highly efficient and convenient once they were familiar with it. P5 and P9 further specified that the gaze selection in (G) prevented them from selecting one keyword while looking at another, which is an inherent limitation of gaze selection and required time to adapt. For (T), all participants reported that the slow typing speed and typos led to high cognitive load and even frustration. P1, P6 and P10 further commented that the need to structure the notes, even with minimal keywords, greatly increased their cognitive load. These contribute to significantly higher distraction, cognitive load, and frustration in (T).
In terms of subtlety, 6 participants (P2, P4, P9, P10, P11, P12) agreed that (G) only requiring minimal eye and finger movement made it imperceptible. On the contrary, 7 participants (P1, P2, P4, P6, P10, P11, P12) pointed out that looking down at the smartphone was quite noticeable and tiring in both (T) and (S), especially in (T) due to the longer duration for manual typing. These result in (G) being the most subtle method with the least physical effort. For social acceptance, most participants commented that taking notes during a speech was common and would not be considered impolite, even if constantly looking at the smartphone. This causes no statistical significance in social acceptance. For usability, all participants appreciated the context keyword feature in (G) as automatic highlights of the speeches, which was not achieved in (S) due to the need to look down the smartphone. Furthermore, P4, P8, P9 and P10 found the selection feature in (S) and (G) instead of manual typing to be very useful. P10 elaborated that “This combination of AR and AI overcomes all the shortcomings of real-time note-taking.” These result in significantly better usability in (G). For note refinement, only P2 and P10 selected two notes recorded with only keywords to review the transcripts but did not modify or refine the notes, respectively. Besides, extractive note-taking, which typically involves directly copying, highlighting or quoting the original text, and abstractive note-taking, which generally involves rephrasing or summarizing based on the original context, are two critical concepts in note-taking. Since sentences of notes are generated by the LLM based on the selected keywords, the GazeNoter system focus on achieving abstractive note-taking for both within-context and beyond-context notes. However, since transcripts are also recorded in note refinement, extractive note-taking is inherently accomplished as well. Based on the results of 4 questions for notes, and with only P2 and P10 reviewing the extractive notes in note refinement, users expressed greater satisfaction with the abstractive notes from GazeNoter compared to the extractive notes.
@@ 5. User Study 2: Walking Meeting
In the previous study, the performance of GazeNoter in the speech scenario was evaluated in a static condition with less interference. We further intended to evaluate its performance in a mobile condition, so attending walking meetings, an emerging use case and research area (Damen et al., 2020a, b; Haliburton et al., 2023), was the scenario examined in this study. Moreover, walking meetings involve interactive communication and occasional face-to-face discussion, resulting in frequent eye contact and less time for note-taking. Hence, the mental pressure in this scenario could be higher compared to simply listening to speeches in the previous study.
@@
Study 2 setup for walking meetings in an area measuring 3.5m $\times$ 4m with an introducer.
@@ 5.1. Setup, Task and Procedure
The setup was similar to the previous study. 12 participants (6 female) aged 22-27 (mean: 25.08), who had not participated in the prior study, were recruited for this study. Participants were compensated with 10 USD for their time. Instead of sitting on a chair, an introducer and a participant walked around a room measuring 3.5m $\times$ 4m for walking meetings. The introducer initiated a topic and spent 3 to 5 minutes describing it, followed by a discussion with the participant during walking meetings. Each walking meeting was about 10 minutes. The participant could freely take notes throughout the description and discussion. However, the participant might need to review the notes for discussion with the introducer. This increased the mental pressure and required more urgent real-time note-taking. As in the previous study, three methods (T), (S) and (G) were examined in each walking meeting, and an additional comparison between (A) and (G) was performed. Similarly, the order of the methods was counterbalanced, and participants filled out a questionnaire and were interviewed after the experiment. The study took about two hours.
@@ 5.2. Results and Discussion
Study 2 quantitative results of the three methods, (T), (S) and (G). Showing (G) outperforms (S) and (T) in note quantitative and quality. Furthermore, participants were willing to spend more time taking notes in dynamic conditions when using (G). The dashed line represents the qualitative results of the additional comparison between (A) and (G), illustrating that note outcome of (G) excels in terms of matching user intention and serves as a more effective reminder.
The results are shown in Figure 10 to 11. Repeated measures ANOVA and Bonferroni correction for post-hoc pairwise tests were used to analyze the objective, quantitative data. A Friedman test and Wilcoxon signed-rank tests with Bonferroni correction for post-hoc pairwise tests were utilized to analyze the subjective questionnaire.
@@ 5.2.1. Quantitative Results
For the objective and quantitative results, significant main effects are revealed in number of notes ($F_{2,22}=9.81,p<0.01$), time of a note ($F_{2,22}=26,p<0.001$) and number of keywords per note ($F_{2,22}=16.97,p<0.001$). Post-hoc pairwise tests show significant differences between (T, G) and (S, G) in the number of notes and time of a note, and among all pairs in the number of keywords per note. The percentage of notes with complete sentences is 23% in (T), 37% in (S), and 63% in (G). The quick note percentage is 42% in (S) and 26% in (G) as shown in the note type distribution of Figure 16 in the appendix. The average time of a quick note in (S) is 5.96s for context keywords only and 7.45s for customized keywords included. The average time of a quick note in (G) is 4.76s for context keywords only and 6.26s for customized keywords included. This scenario leads participants to spend less time on note-taking and prefer quick notes compared to the previous study. For the results of the subjective questionnaire, The beyond-context note percentage is 14% in (S) and 18% in (G) (Figure 16). The derivative keywords are displayed 13% of the total note-taking time in (G).
5.2.2. Subjective Scale Results
For the results of the subjective questionnaire, significant main effects are revealed in intention ($\chi^{2}(2)=22.14,p<0.001$), reminding ($\chi^{2}(2)=15.95,p<0.001$), quality ($\chi^{2}(2)=18.57,p<0.001$), inspiration ($\chi^{2}(2)=13.35,p<0.01$), distraction ($\chi^{2}(2)=22.37,p<0.001$), cognitive load ($\chi^{2}(2)=14.63,p<0.01$), frustration ($\chi^{2}(2)=14.00,p<0.01$), physical effort ($\chi^{2}(2)=17.15,p<0.001$), subtlety ($\chi^{2}(2)=22.14,p<0.001$), social acceptance ($\chi^{2}(2)=17.66,p<0.01$), and usability ($\chi^{2}(2)=21.14,p<0.001$). Post-hoc pairwise tests show significant differences among all pairs in all factors, except between (S, T) in social acceptance. The additional subjective scales of the 4 questions for each note are shown in Figure 19. The differences among the methods are more apparent in this scenario. For the additional comparison between (A) and (G), significant main effects are revealed in intention ($\chi^{2}(1)=12.00,p<0.01$) and reminding ($\chi^{2}(1)=12.00,p<0.01$), and no significant main effects are found in quality ($\chi^{2}(1)=0.14,p=0.70$) and inspiration ($\chi^{2}(1)=2.67,p=0.10$), consistent with the previous study.
@@ 5.2.2. Subjective Scale Results
Study 2 qualitative and subjective scale of the user study on a 7-point Likert scale. The dashed line on the left highlights the 4 questions for notes from each method. The dashed line on the right highlights the scales that lower scores indicate better performance. Results show that (G) significantly outperforms (S) and (T) in various accepts, including intention, reminding, quality, inspiration, distraction, cognitive load, frustration, physical effort, subtlety, and usability.